The income elasticity of restaurant meals is 0.6. If customers income becomes 300,000 LBP instead of 200,000 LBP , the quantity demanded will
1
Expert's answer
2020-04-30T09:53:46-0400

"E"dI"=\frac{\Delta Q\%}{\Delta I\%};"

"\Delta I\%=\frac {300000}{200000}×100\%-100\%=50\%;"

"0.6=\frac {\Delta Q}{50\%};"

"\Delta Q=0.6×50\%=30\%;"

Answer: the quantity demanded  will be increased by 30%.